Tanzanian Excursions generates revenue of $7,000 per person on its five-day package tours to wildlife parks in Tanzania. The variable costs per person are as follows:

Airfare

$        1,900

Hotel accommodations

           2,900

Meals

              450

Ground transportation

              350

Park tickets and other costs

              600

Total

$        6,200

Each tour is for one person, so the above revenue and variable costs per person are effectively per package tour. Annual fixed costs include leased office space, management and support staff salaries, advertising, utilities, and equipment rental. These fixed costs total $540,000 for the year.

  1. Calculate the number of package tours that must be sold to break-even.
  1. Herb Kirkstreet, the owner-manager of Tanzanian Excursions, wants target operating income of $100,000 for the year, which he hopes to use to buy a Tesla Model S for his personal use. Calculate a) the number of package tours, and b) the total revenue needed to generate a target operating income of $100,000.

Herb Kirkstreet is considering two alternative strategies to increase profitability by reducing costs, as outlined in questions 3 and 4 below. Evaluate the alternatives in these two questions independent of each another. For both questions 3 and 4, focus on the short-term effects (i.e., the following year). Assume that the proposed changes will not have any effect on the number of tours for the following year, which is 800 tours at a price per tour of $7,000.

  1. Tanzanian Excursions has been contracting with the Ritz-Carlton to provide luxury hotel accommodations and ground transportation. Kirkstreet is considering moving the contract to a lower cost vendor for hotel accommodations and ground transportation. This change would reduce the variable hotel and ground transportation costs per tour shown above by 10% and 30%, respectively. The new, lower-cost vendor would also take over some of the registration and record keeping functions currently performed by Tanzanian Excursions staff and this would allow the company to eliminate one of the lower level staff positions, reducing the fixed cost per year by $60,000.
    1. what would be the total reduction in cost from this change at the assumed volume of 800 tours?
    2. what would be the total operating income at the assumed volume of 800 tours?
  1. Alternative 2 is to substantially decrease advertising expenditures, which would decrease total fixed costs by 20%. If this alternative is adopted,
    1. what would be the total operating income at the assumed volume of 800 tours?
    2. what is the difference between the breakeven number of tours without this change in advertising expenditures and the breakeven number of tours with the change in advertising expenditures?
  1. For this question, focus on longer-term effects of changes in the cost structure. Because tours are booked about one year in advance, changes in cost structure made in the current year might affect sales two or more years down the road. What longer-term effects would you want Herb Kirkstreet to consider if he implements the cost reductions? List one consideration for each alternative proposed in questions 3 and 4.
Answer #1

1. Number of tours to be sold in order to break even = Fixed Cost / Contribution Margin per Tour = $ 540,000 / $ ( 7,000 - 6,200 ) = 675 tours.

2. a. Number of tours needed to earn operating income of $ 100,000 = $ ( 540,000 + 100,000 ) / $ 800 = 800 tours.

b. Total revenue to be generated to earn $ 100,000 = ( Fixed Costs + Target Operating Income) / Contribution margin Ratio = ( $ [( 540,000 + 100,000 ) / $ 800] * 7,000 = $ 5,600,000.

3. a. Total reduction in cost = ($ 2,900 x 10 % + $ 350 x 30 %) * 800 + $ 60,000 = $ 376,000.

b. Total operating income at the assumed volume of 800 package tours = $ 100,000 + $ 376,000 = $ 476,000.

4. a. Saving in fixed costs = $ 540,000 x 20 % = $ 108,000.

Total operating income at an assumed volume of 800 tours = $ 100,000 + $ 108,000 = $ 208,000.

New break-even number of units = $ ( 540,000 - 108,000 ) / $ 800 = 540 tours.

Therefore, difference in break-even number of tours = 675 - 540 = 135 tours.

5. Longer term effects:

  • The hotels might not be willing to offer the concessions in accommodation or ground transportation for very long.
  • A cutback in advertising spends might begin to have an adverse impact on the tour revenues beyond the short to medium term.